OpenStudy (anonymous):

What is the volume of a triangular prism with a height of 3 centimeters and a base that is an equilateral triangle with sides of 4 centimeters? Answer 24 cm3 12sqr 3cm3 48 cm3 8sqr 3cm3

OpenStudy (lgbasallote):

volume = base area x height base area = (sq 3)/4 (l^2) area = sq 3 /4 (16) area = 4 (sq 3) volume = 4(sq 3) x 3 volume = 12 sq 3 cm^3

OpenStudy (anonymous):

thank you : )